A Sample of gas (1764) at 6.00c and orah Pressure is heated to 2540 and 2.45am farol pressure. what is le fall some
Answer #1

P = T = 0.92 6°C V = 1076L OR HB (273+6) k = 279 k. = lly, 2015 atm V2 = ? pe = T2 = 254° c = ( 254 +273) K = 527 K • PV = nRHope you will like my answer.